# Multiplication of 0-1 matrices via clustering

### v2

@@ Abstract
We study applications of clustering (in particular the $k$-center clustering problem) in the design of efficient and practical deterministic algorithms for computing an approximate and the exact arithmetic matrix product of two 0-1 rectangular matrices $A$ and $B$ with clustered rows or columns, respectively. Let $\lambda_{A}$ and $\lambda_{B}$ denote the minimum maximum radius of a cluster in an $\ell$-center clustering of the rows of $A$ and in a $k$-center clustering of the columns of $B,$ respectively. In particular, when $A$ and $B$ are square matrices of size $n\times n$, we obtain the following results.
  1. 1.
    A simple deterministic algorithm that approximates each entry of the arithmetic matrix product of $A$ and $B$ within an additive error of at most $2\lambda_{A}$ in $O(n^{2}\ell)$ time or at most $2\lambda_{B}$ in $O(n^{2}k)$ time.
  2. 2.
    A simple deterministic preprocessing of the matrices $A$ and $B$ in $O(n^{2}\ell)$ time or $O(n^{2}k)$ time after which every query asking for the exact value of an arbitrary entry of the arithmetic matrix product of $A$ and $B$ can be answered in $O(\lambda_{A})$ time or $O(\lambda_{B})$ time, respectively.
  3. 3.
    A simple deterministic algorithm for the exact arithmetic matrix product of $A$ and $B$ running in time $O(n^{2}(\ell+k+\min\{\lambda_{A},\lambda_{B}\}))$.
Keywords:
a
@@ 1 Introduction
The arithmetic matrix product of two 0-1 matrices is closely related to the Boolean one of the corresponding Boolean matrices. For square $n\times n$ matrices, both can be computed in $O(n^{2.372})$ time [1, 23]. Both are basic tools in science and engineering. Unfortunately, no truly subcubic practical algorithms for any of them are known. Therefore, many researchers studied the complexity of these products for special input matrices, e.g., sparse or structured matrices [3, 4, 11, 15, 21, 24], providing faster and often more practical algorithms.
@@
The method of multiplying matrices with clustered rows or columns, proposed for Boolean matrix product in [4] and subsequently generalized in [11, 15] and used in [2], relies on the construction of an approximate spanning tree of the rows of the first input matrix or the columns of the second input matrix in a Hamming space. Then, each column or each row of the product matrix is computed with the help of a traversal of the tree in time proportional to the total Hamming cost of the tree up to a logarithmic factor. Simply, the next entry in a column or a row in the product matrix can be obtained from the previous one in time roughly proportional to the Hamming distance between the consecutive (in the tree traversal) corresponding rows or columns of the first or the second input matrix, respectively. Thus, in case the entire tree cost is substantially subquadratic in $n,$ the total running time of this method becomes substantially subcubic provided that a good approximation of a minimum spanning tree of the rows of the first input matrix or the columns of the second one can be constructed in a substantially subcubic time. As for simplicity and practicality, a weak point of this method is that in order to construct such an approximation relatively quickly, it employs a randomized dimension reduction.
@@
In case of the arithmetic matrix product of 0-1 matrices, in some cases, a faster approximate arithmetic matrix multiplication can be more useful [8, 21]. Among other things, it can enable to identify largest entries in the product matrix and it can be also used to provide a fast estimation of the number of: the so called witnesses for the Boolean product of two Boolean matrices [14], triangles in a graph, or more generally, subgraphs isomorphic to a small pattern graph [12] etc. There is a number of results on approximate arithmetic matrix multiplication, where the quality of approximation is expressed in terms of the Frobenius matrix norm $||\ ||_{F}$ (i.e., the square root of the sum of the squares of the entries of the matrix) [8, 21].
Cohen and Lewis [8] and Drineas et al. [9] used random sampling to approximate arithmetic matrix product. Their papers provide an approximation $D$ of the matrix product $AB$ of two $n\times n$ matrices $A$ and $B$ such that $||AB-D||_{F}=O(||AB||_{F}/\sqrt{c})$, for a parameter $c>1$ (see also [21]). The approximation algorithm in [9] runs in $O(n^{2}c)$ time. Drineas et al. [9] also derived bounds on the entrywise differences between the exact matrix product and its approximation. Unfortunately, the best of these bounds is $\Omega(M^{2}n/\sqrt{c})$, where $M$ is the maximum value of an entry in $A$ and $B.$ By using a sketch technique, Sarlós [22] obtained the same Frobenius norm guarantees, also in $O(n^{2}c)$ time. However, he derived stronger individual upper bounds on the additive error of each entry $D_{ij}$ of the approximation matrix $D$. They are of the form $O(||A_{i*}||_{2}||B_{*j}||_{2}/\sqrt{c})$, where $A_{i*}$ and $B_{*j}$ stands for the $i$-th row of $A$ and the $j$-th column of $B,$ respectively, that hold with high probability. More recently, Pagh [21] presented a randomized approximation $\tilde{O}(n(n+c))$-time algorithm for the arithmetic product of $n\times n$ matrices $A$ and $B$ such that each entry of the approximate matrix product differs at most by $||AB||_{F}/\sqrt{c}$ from the correct one. His algorithm first compresses the matrix product to a product of two polynomials and then uses the fast Fourier transform to multiply the polynomials. Subsequently, Kutzkov [20] developed analogous deterministic algorithms employing different techniques. For approximation results related to sparse arithmetic matrix products, see [19, 21].
@@ 1.1 Our contributions
In this paper, we exploit the possibility of applying the classic simple
$2$-approximation algorithm for the $k$ center clustering problem [16] in order to derive efficient and practical deterministic algorithms for computing an approximate and the exact arithmetic matrix product of two 0-1 rectangular matrices $A$ and $B$ with clustered rows or columns, respectively.
The $k$-center clustering problem in a Hamming space $\{0,1\}^{d}$ is for a set $P$ of $n$ points in $\{0,1\}^{d}$ to find a set $T$ of $k$ points in $\{0,1\}^{d}$ that minimize $\max_{v\in P}\min_{u\in T}\mathrm{ham}(v,u),$ where $\mathrm{ham}(v,u)$ stands for the Hamming distance between $v$ and $u,$ (i.e., the number of coordinate positions they differ from each other). Each center in $T$ induces a cluster consisting of all points in $P$ for which it is the nearest center.
Let $\lambda_{A}$ and $\lambda_{B}$ denote the minimum maximum radius of a cluster in an $\ell$-center clustering of the rows of $A$ or in a $k$-center clustering of the columns of $B,$ respectively. Assuming that $A$ and $B$ are of sizes $p\times q$ and $q\times r,$ respectively, we obtain the following results.
@@
    A simple deterministic algorithm that approximates each entry of the arithmetic matrix product of $A$ and $B$ within an additive error of at most $2\lambda_{A}$ in $O(pq\ell+pr)$ time if $p\geq r$ or at most $2\lambda_{B}$ in $O(qrk+pr)$ time if $p\leq r$.
@@
    A simple deterministic preprocessing of the matrices $A$ and $B$ in $O(pq\ell+pr)$ time if $p\geq r$ or $O(qrk+pr)$ time if $p\leq r$ after which every query asking for the exact value of an arbitrary entry of the arithmetic matrix product of $A$ and $B$ can be answered in $O(\lambda_{A})$ time if $p\geq r$ or $O(\lambda_{B})$ time if $p\leq r.$
@@
    A simple deterministic algorithm for the exact arithmetic matrix product of $A$ and $B$ running in time $O(pq\ell+rqk+\min\{pr\lambda_{A}+rq\ell,pr\lambda_{B}+pqk\}).$
@@ 1.2 Techniques
All our main results rely on the classical, simple $2$-approximation algorithm for the $k$-center clustering problem (farthest-point clustering) due to Gonzalez [16] (see also Fact 1). Two of them rely also on the idea of updating the inner product of two vectors $a$ and $b$ in $\{0,1\}^{q}$ over the Boolean or an arithmetic semi-ring to that of two vectors $a^{\prime}$ and $b^{\prime}$ in $\{0,1\}^{q}$, where $a=a^{\prime}$ or $b=b^{\prime},$ in time roughly proportional to $\mathrm{ham}(a,a^{\prime})+\mathrm{ham}(b,b^{\prime})$. The idea has been used in [4, 11, 15]. As in the aforementioned papers, we combine it with a traversal of an approximate minimum spanning tree of the rows or columns of an input matrix in the Hamming space $\{0,1\}^{q}$, where $q$ is the length of the rows or columns (see also Lemma 6).
1.3 Paper organization
The next section contains basic definitions. Section 3 presents our approximation algorithm for the arithmetic product of two 0-1 matrices and the preprocessing enabling efficient answers to queries asking for the value of an arbitrary entry of the arithmetic product matrix. Section 4 is devoted to our algorithm for the exact arithmetic matrix product of two 0-1 matrices. We conclude with a short discussion on possible extensions of our results.
@@ 2 Preliminaries
The transpose of a matrix $D$ is denoted by $D^{\top}.$ If the entries of $D$ are in $\{0,1\}$ then $D$ is a 0-1 matrix.
The Hamming distance between two points $a,\ b$ (vectors) in $\{0,1\}^{d}$ is the number of the coordinates in which the two points differ. Alternatively, it can be defined as the distance between $a$ and $b$ in the $L_{1}$ metric over $\{0,1\}^{d}.$ It is denoted by $\mathrm{ham}(a,b).$
The $k$-center clustering problem in a Hamming space $\{0,1\}^{d}$ is as follows: given a set $P$ of $n$ points in $\{0,1\}^{d}$, find a set $T$ of $k$ points in $\{0,1\}^{d}$ that minimize $\max_{v\in P}\min_{u\in T}\mathrm{ham}(v,u).$
The minimum-diameter $k$-clustering problem in a Hamming space $\{0,1\}^{d}$ is as follows: given a set $P$ of $n$ points in $\{0,1\}^{d}$, find a partition of $P$ into $k$ subsets $P_{1},P_{2},\dots,P_{k}$ that minimize $\max_{i\in[k]}\max_{v,u\in P_{i}}\mathrm{ham}(v,u).$ Note that the $k$-center clustering problem could be also termed as the minimum-radius $k$-clustering problem. It is known to be NP-hard and even NP-hard to approximate within $2-\epsilon$ for any constant $\epsilon>0$ [10, 13].
@@
[16] Let $P$ be a set of $n$ points in $\{0,1\}^{d}$, and let $k\in[n].$ There is a simple deterministic $2$-approximation algorithm for the $k$-center clustering and minimum-diameter $k$-clustering problems running in $O(ndk)$ time.
@@ 3 An approximate arithmetic matrix product of 0-1 matrices
procedure $APPROXMMCLUS(A,B,\ell)$
Input: Two 0-1 matrices $A$ and $B$ of sizes $p\times q$ and $q\times r,$ respectively, where $p\geq r,$ and a positive integer $\ell$ not exceeding $p$.
Output: A $p\times r$ matrix $D$, where for $1\leq i\leq p$ and $1\leq j\leq r,$ $D_{ij}$ is an approximation of the inner product $C_{ij}$ of the $i$-th row $A_{i*}$ of $A$ and the $j$-th column $B_{*j}$ of $B.$
  1. 1.
    Determine an approximate $\ell$-center clustering of the rows of the matrix $A$ in $\{0,1\}^{q}.$ For each row $A_{i*}$ of $A$, set $cen_{\ell}(A_{i*})$ to the center of the cluster in the $\ell$-clustering to which $A_{i*}$ belongs.
  2. 2.
    Form the ×$\ell\times q$ matrix $A^{\prime},$ where the $i^{\prime}$-th row is the $i^{\prime}$-th center in the approximate $\ell$-center clustering of the rows of $A.$
  3. 3.
    Compute the arithmetic ×$\ell\times r$ matrix product $C^{\prime}$ of $A^{\prime}$ and $B.$
  4. 4.
    For $1\leq i\leq p$ and $1\leq j\leq r,$ set $D_{ij}$ to $C^{\prime}_{i^{\prime}j}$, where the $i^{\prime}$-th row $A^{\prime}_{i^{\prime}*}$ of $A^{\prime}$ is $cen_{\ell}(A_{i*}).$
For a 0-1 $p\times q$ matrix $A$, let $\lambda(A,\ell,row)$ be the minimum, over all $\ell$-center clusterings of the rows of $A$ in the Hamming space $\{0,1\}^{q}$, of the maximum Hamming distance between a center of a cluster and a member of the cluster. Similarly, for a 0-1 $q\times r$ matrix $B,$ let $\lambda(B,k,col)$ be the minimum, over all $k$-center clusterings of the columns of $B$ in the Hamming space $\{0,1\}^{q}$, of the maximum Hamming distance between a center of a cluster and a member of the cluster.
@@
Suppose a $2$-approximation algorithm for the $\ell$-center clustering is used in $APPROXMMCLUS(A,B,\ell)$ and $C$ stands for the arithmetic product of $A$ and $B.$ Then, for $1\leq i\leq p$ and $1\leq j\leq r,$ $|C_{ij}-D_{ij}|\leq 2\lambda(A,\ell,row)$.
@@
Recall that $p\geq r$ is assumed in the input to $APPROXMMCLUS(A,B,\ell)$. For $1\leq i\leq p$ and $1\leq j\leq r,$ $D_{ij}$ is the inner product of $cen_{\ell}(A_{i*})$, where $\mathrm{ham}(A_{i*},cen_{\ell}(A_{i*}))\leq 2\lambda(A,\ell,row)$, with $B_{*j}.$ Hence, $C_{ij}$, which is the inner product of $A_{i*}$ with $B_{*j},$ can differ at most by $2\lambda(A,\ell,row)$ from $D_{ij}.$
@@
$APPROXMMCLUS(A,B,\ell)$ can be implemented in
$O(pq\ell+pr+T(\ell,q,r))$ time.
@@
Recall that $p\geq r.$ Step 1, which includes the assignment of the closest center to each row of $A$, can be done in $O(pq\ell)$ time by using Fact 1, i.e., the classic algorithm of Gonzalez [16]. Step 2 takes $O(\ell q)$ time, which is $O(T(\ell,q,r))$ time. Finally, Step 3 takes $T(\ell,q,r)$ time while Step 4 can be done in $O(pr)$ time. Thus, the overall time is $O(pq\ell+pr+T(\ell,q,r)).$
We can use the straightforward $O(sqt)$-time algorithm for the multiplication of two matrices of sizes $s\times q$ and $q\times t,$ respectively. Since $T(\ell,q,r)=O(\ell qr)=O(\ell qp)$ if $p\geq r,$ Lemmata 1 and 2 yield the first part (1) of our first main result (Theorem 3.1 below), for $p\geq r.$ Its second part (2) for $p\leq r$ follows from the first part by $(AB)^{\top}=B^{\top}A^{\top}.$ Note that then the number of rows in $B^{\top}$, which is $r$, is not less than the number of columns in $A^{\top}$, which is $p$. Simply, we run $APPROXMMCLUS(B^{\top},A^{\top},k)$ in order to compute an approximation of the transpose of the arithmetic matrix product of $A$ and $B.$ Note also that a $k$-clustering of columns of $B$ is equivalent to a $k$-clustering of the rows of $B^{\top}$ and that $\lambda(B^{\top},k,row)=\lambda(B,k,col).$

### v3

# Jesper Jansson    Mirosław Kowaluk    Andrzej Lingas    Mia Persson

@@ Abstract
We study applications of clustering (in particular, the $k$-center clustering problem) in the design of efficient and practical algorithms for computing an approximate and the exact arithmetic matrix product of two 0-1 rectangular matrices with clustered rows or columns, respectively. Our results in part can be regarded as an extension of the clustering-based approach to Boolean square matrix multiplication due to Arslan and Chidri (CSC 2011). First, we provide a simple and efficient deterministic algorithm for approximate matrix product of 0-1 matrices, where the additive error is proportional to the minimum maximum radius in an $\ell$-center clustering of the rows of the first matrix or an $k$-center clustering of the columns of the second matrix. Next, we use the approximation algorithm as a preprocessing after which a query asking for the exact value of an arbitrary entry in the product matrix can be answered in time proportional to the additive error. As a consequence, we obtain a simple deterministic algorithm for the exact matrix product of 0-1 matrices. We also present an improved simple deterministic algorithm for the exact product and in addition, faster analogous randomized algorithms for an approximate and the exact matrix products of 0-1 matrices based on randomized $\ell$ and $k$-center clustering.
@@ 1 Introduction
The arithmetic matrix product of two 0-1 matrices is closely related to the Boolean one of the corresponding Boolean matrices. Both are basic tools in science and engineering. For square $n\times n$ matrices, both can be computed in $O(n^{2.372})$ time by using fast matrix multiplication algorithms based on algebra [ADV25, VXZ24]. Unfortunately, the latter algorithms suffer from huge overheads and no truly subcubic practical algorithms for any of these two problems is known. Therefore, many researchers studied the complexity of these products for special input matrices, e.g., sparse or structured matrices [AC10, Emi24, BL, FJLL18, GL03, P13, YZ], providing faster and often more practical algorithms.
@@
The method of multiplying matrices with clustered rows or columns, proposed for Boolean matrix product in [BL] and subsequently generalized in [FJLL18, GL03] and used in [AMBFGR_25], relies on the construction of an approximate minimum spanning tree of the rows of the first input matrix or the columns of the second input matrix in a Hamming space. Then, each column or each row of the product matrix is computed with the help of a traversal of the tree in time proportional to the total Hamming cost of the tree up to a logarithmic factor. Simply, the next entry in a column or a row in the product matrix can be obtained from the previous one in time roughly proportional to the Hamming distance between the consecutive (in the tree traversal) corresponding rows or columns of the first or the second input matrix, respectively. Thus, in case the entire tree cost is substantially subquadratic in $n,$ the total running time of this method becomes substantially subcubic provided that a good approximation of a minimum spanning tree of the rows of the first input matrix or the columns of the second one can be constructed in substantially subcubic time. As for simplicity and practicality, a weak point of this method is that in order to construct such an approximation relatively quickly, it employs a rather involved randomized algorithm.
Arslan and Chidri presented a simple algorithm for the Boolean matrix product of two input Boolean matrices with rows (in the first matrix) and columns (in the other matrix) within a small diameter $d$ (in terms of the Hamming distance) in [AC11]. The basic idea is to compute the inner Boolean product of a representative of the rows of the first matrix and a representative of columns of the second matrix and then relatively cheaply recover the entries of the Boolean product matrix from the inner product. Their algorithm runs in $O(dn^{2})$ time for $n\times n$ matrices. They also presented a generalization of their algorithm to include grouping the rows of the first matrix as well as the columns of the second matrix into $k$ clusters of maximum radius within $2$ of the minimum $s.$ They incorporate Gonzalez’s $2$-approximation algorithm for the $k$-center problem [Gon85] to achieve such a $k$-clustering. Their general algorithm runs in $O((k+s)n^{2})$ time [^1] They also claim that one can speed up their general algorithm by using the fast $2$-approximation algorithm for the $k$-center problem due to Feder and Greene [FG88] instead of Gonzalez’s one. Unfortunately, the former algorithm seems to have rather an exponential hidden dependence on the dimension (in this case $n$) in view of Lemmata 4.2 and 4.3 in [FG88] than a linear one as it is assumed in [AC11]..
@@
In case of the arithmetic matrix product of 0-1 matrices, in some cases, a faster approximate arithmetic matrix multiplication can be more useful [CL99, P13]. Among other things, it can enable to identify largest entries in the product matrix and it can be also used to provide a fast estimation of the number of: the so called witnesses for the Boolean product of two Boolean matrices [GKL09], triangles in a graph, or more generally, subgraphs isomorphic to a small pattern graph [FKL15] etc. There is a number of results on approximate arithmetic matrix multiplication, where the quality of approximation is expressed in terms of the Frobenius matrix norm $||\ ||_{F}$ (i.e., the square root of the sum of the squares of the entries of the matrix) [CL99, P13].
[…]
@@ 1.1 Our contributions
In this article, we exploit the possibility of applying the classic simple $2$-approximation algorithm for the $k$ center clustering problem [Gon85] in order to derive efficient and practical deterministic algorithms for computing an approximate and the exact arithmetic matrix product of two 0-1 rectangular matrices $A$ and $B$ with clustered rows or columns, respectively. In addition, we consider also the possibility of applying the recent faster randomized algorithm for $k$ center clustering from [KLP26] instead of that from [Gon85].
Our results are based on the idea of using cluster representatives/centers to compute faster a smaller matrix product and then relatively cheaply updating the small product to the target one. This natural idea has been used already in [AC11] [^2] We came across this idea independently without knowing [AC11] before writing the preliminary conference version of this article.. There are however several differences between our approach and that in [AC11]. We consider the more general problem of computing the arithmetic matrix product of 0-1 rectangular matrices while [AC11] is concerned with the Boolean matrix product of square Boolean matrices. We provide also cluster-based approximation and preprocessing for a query on the value of single entry of the aforementioned arithmetic product. In [AC11], the rows of the first matrix are grouped in the same number of clusters as the columns of the second matrix, while in our case the numbers are in general different. Also, in [AC11] representatives/centers of both row clusters and column clusters are used to compute the smaller product while we use in the basic deterministic version either the row ones or the column ones.
The $k$-center clustering problem in a Hamming space $\{0,1\}^{d}$ is for a set $P$ of $n$ points in $\{0,1\}^{d}$ to find a set $T$ of $k$ points in $\{0,1\}^{d}$ that minimize $\max_{v\in P}\min_{u\in T}\mathrm{ham}(v,u),$ where $\mathrm{ham}(v,u)$ stands for the Hamming distance between $v$ and $u$ (i.e., the number of coordinates where $v$ and $u$ differ). Each center in $T$ induces a cluster consisting of all points in $P$ for which it is the nearest center.
Let $\lambda(A,\ell,row)$ and $\lambda(B,k,col)$ denote the minimum maximum Hamming distance between a row of $A$ or a column of $B$ and the closest center in an $\ell$-center clustering of the rows of $A$ or in a $k$-center clustering of the columns of $B,$ respectively. In particular, when $A$ and $B$ are square matrices of size $n\times n$, we obtain the following results.
@@
    A simple deterministic algorithm that approximates each entry of the arithmetic matrix product of $A$ and $B$ within an additive error of at most $2\lambda(A,\ell,row)$ in $O(n^{2}\ell)$ time or at most $2\lambda(B,k,col)$ in $O(n^{2}k)$ time. Similarly, a simple randomized algorithm that for any $\epsilon\in(0,\frac{1}{2})$ approximates each entry of the product within an additive error of with high probability (w.h.p.) at most $(2+\epsilon)(\lambda(A,\ell,row)+\lambda(B,k,col))$ in time $O(n\log n(n+\ell+k)/\epsilon^{2}+\ell nk)$.
@@
    A simple deterministic preprocessing of the matrices $A$ and $B$ in $O(n^{2}\ell)$ time or $O(n^{2}k)$ time after which every query asking for the exact value of an arbitrary entry of the arithmetic matrix product of $A$ and $B$ can be answered in $O(\lambda(A,\ell,row))$ time or $O(\lambda(B,k,col))$ time, respectively. Similarly, a simple randomized preprocessing of the matrices in time $O(n\log n(n+\ell+k)+\ell nk)$ after which every query asking for the exact value of an arbitrary entry of the product of the matrices can be answered in $O(\lambda(A,\ell,row)+\lambda(B,k,col))$ time w.h.p.
@@
    A simple deterministic algorithm for the exact arithmetic matrix product of $A$ and $B$ running in time $O(n^{2}(\ell+k+\min\{\lambda(A,\ell,row),\lambda(B,k,col)\}))$. Alternatively, a simple randomized algorithm for the exact product running in time $O(n\log n(n+\ell+k)+\ell nk+n^{2}(\lambda(A,\ell,row)+\lambda(B,k,col))$ w.h.p.
@@ 1.2 Techniques
All our main deterministic results rely on the classical, simple $2$-approximation algorithm for the $k$-center clustering problem (farthest-point clustering) due to Gonzalez [Gon85], whose properties are summarized in Fact 1 below. Our results also rely on the idea of updating the inner product of two vectors $a$ and $b$ in $\{0,1\}^{q}$ over the Boolean or an arithmetic semi-ring to that of two vectors $a^{\prime}$ and $b^{\prime}$ in $\{0,1\}^{q}$, in time roughly proportional to $\mathrm{ham}(a,a^{\prime})+\mathrm{ham}(b,b^{\prime})$. The idea has been used in [AC11, BL, FJLL18, GL03]. As in some of the aforementioned articles, we combine it with a traversal of an approximate minimum spanning tree of the rows or columns of an input matrix in the Hamming space $\{0,1\}^{q}$, where $q$ is the length of the rows or columns (see Lemma 8).
1.3 Article organization
The next section contains basic definitions. Section 3 presents our approximation algorithms for the arithmetic product of two 0-1 matrices and the preprocessing enabling efficient answers to queries asking for the value of an arbitrary entry of the arithmetic product matrix. Section 4 is devoted to our algorithms for the exact arithmetic matrix product of two 0-1 matrices. We conclude with a short discussion on potential extensions of our results.
@@ 2 Preliminaries
For two vectors $(a_{1},\dots,a_{d})$ and $(b_{1},\dots,b_{d})$ in $\mathbb{R}^{d},$ their inner product is equal to $\sum_{\ell=1}^{d}a_{\ell}b_{\ell}.$ The transpose of a matrix $D$ is denoted by $D^{\top}.$ If the entries of $D$ are in $\{0,1\}$ then $D$ is a 0-1 matrix.
The Hamming distance between two points a,ba,\ b (vectors) in $\{0,1\}^{d}$ is the number of coordinates in which the two points differ. Alternatively, it can be defined as the distance between $a$ and $b$ in the $L_{1}$ metric over $\{0,1\}^{d}.$ It is denoted by $\mathrm{ham}(a,b).$
An event is said to hold with high probability (w.h.p. for short) in terms of a parameter $N$ related to the input size if it holds with probability at least $1-\frac{1}{N^{\alpha}},$ for any constant $\alpha$ not less than $1.$
The $k$-center clustering problem in a Hamming space is as follows: Given a set $P$ of $n$ points in a Hamming space $\{0,1\}^{d}$, find a set $T$ of $k$ points in $\{0,1\}^{d}$ that minimize $\max_{v\in P}\min_{u\in T}\mathrm{ham}(v,u).$ It is known to be NP-hard already for $k=1$ [FL97] and also NP-hard to approximate within a ratio of $2-\epsilon$ for any constant $\epsilon>0$ when $k$ is part of the input [FG88, GJL04]. Note that according to the problem definition, the points in $T$ do not have to belong to the set $P$. The minimum-diameter $k$-clustering problem in a Hamming space is: Given a set $P$ of $n$ points in a Hamming space $\{0,1\}^{d}$, find a partition of $P$ into $k$ subsets $P_{1},P_{2},\dots,P_{k}$ that minimize $\max_{i\in[k]}\max_{v,u\in P_{i}}\mathrm{ham}(v,u).$ The $k$-center clustering problem can also be termed the minimum-radius $k$-clustering problem.
@@
[Gon85] (Gonzalez’s algorithm) There is a simple deterministic
22-approximation algorithm for the $k$-center clustering and minimum-diameter $k$-clustering problems in a Hamming space running in $O(ndk)$ time. Moreover, the approximate solution returned by Gonzalez’s algorithm consists entirely of points from the input set $P$.
We will use the last property given in Fact 1 (that is, the approximate solutions found by Gonzalez’s algorithm consist of points from the input only) in the proof of Theorem 4.1.
A combination of a variant of randomized dimension reduction in Euclidean spaces in the spirit of Johnson and Lindenstrauss [JL84] given by Achlioptas in [A03] and the observation that the Hamming distance between two 0-1 vectors is equal to their squared $L_{2}$ distance makes it possible to decrease the dimension to a logarithmic one in the number of input points, at the cost of worsening the approximation guarantee by some arbitrarily small $\epsilon$. Taking into account the time needed to compute the images of the input points in the subspace of logarithmic dimension, this yields the following fact whose full proof can be found in [KLP26].
Fact 2
[KLP26] For any fixed $\epsilon\in(0,1/2)$, there is a randomized algorithm for the $k$-center clustering problem in a Hamming space running in $O(n\log n(d+k)/\epsilon^{2})$ time that w.h.p. computes a $(2+\epsilon)$-approximation of an optimal $k$-center clustering.
@@ 3 An approximate arithmetic matrix product of 0-1 matrices
procedure $MMCLUS$-$Approx(A,B,\ell)$ Input: Two 0-1 matrices $A$ and $B$ of sizes $p\times q$ and $q\times r,$ respectively, where $p\geq r,$ and a positive integer $\ell$ not exceeding $p$. Output: A $p\times r$ matrix $D$, where for $1\leq i\leq p$ and $1\leq j\leq r,$ $D_{ij}$ is an approximation of the inner product $C_{ij}$ of the $i$-th row $A_{i*}$ of $A$ and the $j$-th column $B_{*j}$ of $B.$ 1. Determine an approximate solution to the $\ell$-center clustering of the rows of the matrix $A$ in $\{0,1\}^{q}.$ For each row $A_{i*}$ of $A$, set $cen_{\ell}(A_{i*})$ to its closest center, with ties broken arbitrarily. 2. Form the $\ell\times q$ matrix $A^{\prime},$ where the $i^{\prime}$-th row is the $i^{\prime}$-th center in the approximate $\ell$-center clustering of the rows of $A.$ 3. Compute the arithmetic $\ell\times r$ matrix product $C^{\prime}$ of $A^{\prime}$ and $B.$ 4. For $1\leq i\leq p$ and $1\leq j\leq r,$ set $D_{ij}$ to $C^{\prime}_{i^{\prime}j}$, where the $i^{\prime}$-th row $A^{\prime}_{i^{\prime}*}$ of $A^{\prime}$ is $cen_{\ell}(A_{i*}).$
See Fig. 1 for an illustration of how $C^{\prime}$ is computed from $A^{\prime}$ and $B$ in step 3. After having computed $C^{\prime}$, the algorithm then sets each entry $D_{ij}$ of the approximate matrix $D$ to the entry of $C^{\prime}$ holding the inner product of the center closest to the $i$th row and the $j$th column of $B$ in step 4.
Figure 1: Computing the matrix $C^{\prime}$ in step 3 of $MMCLUS$-$Approx(A,B,\ell)$.
For a 0-1 $p\times q$ matrix $A$, let $\lambda(A,\ell,row)$ be the minimum, over all $\ell$-center clusterings of the rows of $A$ in the Hamming space $\{0,1\}^{q}$, of the maximum Hamming distance between a row of $A$ and the closest center. Similarly, for a 0-1 $q\times r$ matrix $B,$ let $\lambda(B,k,col)$ be the minimum, over all $k$-center clusterings of the columns of $B$ in the Hamming space $\{0,1\}^{q}$, of the maximum Hamming distance between a column of $B$ and the closest center.
@@
Suppose a $2$-approximation algorithm for the $\ell$-center clustering is used in $MMCLUS$-$Approx(A,B,\ell)$ and let $C$ stand for the arithmetic product of $A$ and $B.$ Then, for $1\leq i\leq p$ and $1\leq j\leq r,$ $|C_{ij}-D_{ij}|\leq 2\lambda(A,\ell,row)$.
@@
Recall that $p\geq r$ is assumed in the input to MMCLUSMMCLUS-$Approx(A,B,\ell)$. For $1\leq i\leq p$ and $1\leq j\leq r,$ $D_{ij}$ is the inner product of $cen_{\ell}(A_{i*})$, where $\mathrm{ham}(A_{i*},cen_{\ell}(A_{i*}))\leq 2\lambda(A,\ell,row)$, with $B_{*j}.$ Hence, $C_{ij}$, which is the inner product of $A_{i*}$ with $B_{*j},$ can differ at most by $2\lambda(A,\ell,row)$ from $D_{ij}.$
@@
MMCLUSMMCLUS-$Approx(A,B,\ell)$ can be implemented in $O(pq\ell+pr+T(\ell,q,r))$ time.
@@
Recall that $p\geq r.$ Step 1, which includes the assignment of the closest center to each row of $A$, can be done in $O(pq\ell)$ time by using Fact 1. Step 2 takes $O(\ell q)$ time, which is $O(T(\ell,q,r))$ time. Finally, Step 3 takes $T(\ell,q,r)$ time while Step 4 can be done in $O(pr)$ time. Thus, the overall time is $O(pq\ell+pr+T(\ell,q,r)).$
We can use the straightforward $O(sqt)$-time algorithm for the multiplication of two matrices of sizes $s\times q$ and $q\times t,$ respectively. Since $T(\ell,q,r)=O(\ell qr)=O(\ell qp)$ if $p\geq r,$ Lemmata 1 and 2 yield the first part (1) of our first main result (Theorem 3.1 below), for $p\geq r.$ Its second part (2) for $p\leq r$ follows from the first part by $(AB)^{\top}=B^{\top}A^{\top}.$ Note that then the number of rows in $B^{\top}$, which is $r$, is not less than the number of columns in $A^{\top}$, which is $p$. Simply, we run $MMCLUS$-$Approx(B^{\top},A^{\top},k)$ to compute an approximation of the transpose of the arithmetic matrix product of $A$ and $B.$ Note also that a $k$-clustering of columns of $B$ is equivalent to a $k$-clustering of the rows of $B^{\top}$ and hence $\lambda(B^{\top},k,row)=\lambda(B,k,col).$